OpenStudy (anonymous):

square root of 28 + square root of 20 - square root of 63

OpenStudy (anonymous):

\[\sqrt{28}+\sqrt{20}-\sqrt{63}\]\[=2\sqrt{7}+2\sqrt{5}-3\sqrt{7}=2\sqrt{5}-\sqrt{7}\]This is as much as we can simplify this one.

OpenStudy (saifoo.khan):

thank God you typed that right. lol.

OpenStudy (anonymous):

Whatever. You guys type too fast.

OpenStudy (saifoo.khan):

\[lol.\]

OpenStudy (anonymous):

haha